OpenStudy (anonymous):

Can someone PLEASE help me with this question? I just realized I went about it all wrong! D: (Advanced Algebra with Financial Applications) Choose one of the formulas above. Write a unique real world scenario that would be solved using the formula you chose. Solve for the present value of the scenario you wrote. I chose Formula #1: PV = FV/(1 + i)^(nt) Can someone help come up with a scenario? That's the part I am having the most trouble with so far.

OpenStudy (kewlgeek555):

@AdrianK : Welcome to OpenStudy and I am very sorry one of our trollers had to do this to your post. A moderator will take care of this. :/

OpenStudy (anonymous):

It's okay @kewlgeek555

ganeshie8 (ganeshie8):

try something like below : you want to buy a car in 5 years, the cost of car is... lets say $10,000; For that you're thinking of depositing some money in bank NOW such that in 5 years it grows to $10,000.

ganeshie8 (ganeshie8):

also u need to choose some interest rate also

OpenStudy (anonymous):

Okay, how about 3.5% interest?

OpenStudy (kewlgeek555):

Good luck on your math question! You are in good hands with @ganeshie8 . ;)

ganeshie8 (ganeshie8):

looks good, given that scenario, ur task is to find out HOW MUCH money you need to deposit in bank right NOW

OpenStudy (anonymous):

lol Thank you! @kewlgeek555

OpenStudy (anonymous):

Okay, I think I should be able to do the rest, but would you mind checking my work when I am done?

ganeshie8 (ganeshie8):

good, and yes :)

OpenStudy (anonymous):

Thank you very much (:

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

Will the bank be charging me interest or will I be earning interest? @ganeshie8

OpenStudy (anonymous):

I'll be earning interest right?

ganeshie8 (ganeshie8):

you have deposited ur money into bank so u will be earning interest - ur money will be growing...

OpenStudy (anonymous):

Just double checking! lol

ganeshie8 (ganeshie8):

good to double/triple check :)b

ganeshie8 (ganeshie8):

wat about compounding interval ?

ganeshie8 (ganeshie8):

u going wid annual compounding / monthly ?

OpenStudy (anonymous):

Yes, I am (: I was going to do monthly

ganeshie8 (ganeshie8):

sounds good :)

OpenStudy (anonymous):

Okay, so far I decided to deposite $3,000 (I just guessed to see if it would work) so my formula looks like this: $3,000 = $10,000/(1 + 0.035)^(12*5) (Without doing anything with the $3,000) I ended up with $1,269.34 Is any of this correct?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

nope, u need to use Present Value formula

ganeshie8 (ganeshie8):

I chose Formula #1: PV = FV/(1 + i)^(nt)

OpenStudy (anonymous):

That's the one I used

ganeshie8 (ganeshie8):

your goal is to have $10,000 in 5 years PV gives you how much u need to deposit in bank now

ganeshie8 (ganeshie8):

FV = 10,000 i = 0.0325/12 n = 12 t = 5

OpenStudy (anonymous):

Ohh

ganeshie8 (ganeshie8):

so, \(\large PV = \frac{10,000}{(1+\frac{0.0325}{12})^{12\times 5}}\)

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

where did you get 0.0325?

ganeshie8 (ganeshie8):

oops sorry, it has to be 0.035

ganeshie8 (ganeshie8):

Corrected below : \(\large PV = \frac{10,000}{(1+\frac{0.035}{12})^{12\times 5}}\)

OpenStudy (anonymous):

Okay thats better lol I'll try that one!

ganeshie8 (ganeshie8):

okie

OpenStudy (anonymous):

I got $8,396.70 (rounded off it would be $8,396.71) @ganeshie8

OpenStudy (anonymous):

Ya there? lol @ganeshie8

OpenStudy (anonymous):

I had to get the assignment submitted, so sorry for being impatient. Thank you for all of your help! You are awesome! :D

ganeshie8 (ganeshie8):

hey sorry i just got notification... openstudy is lagging for me :/

ganeshie8 (ganeshie8):

use more words in explaining ur scenario ok ?

ganeshie8 (ganeshie8):

remember, this is a written problem....

ganeshie8 (ganeshie8):

u need to explain ur scenario/example well and also explain how u got ur answer as well

OpenStudy (anonymous):

I did all of that! I should be okay! (: @ganeshie8

ganeshie8 (ganeshie8):

good :)